It is estimated that 4000 of the 10,000 voting
residents of a town are against a new sales tax. If 15
eligible voters are selected at random and asked their
opinion, what is the probability that at most 7 favor
the new tax?
1
Expert's answer
2018-01-19T15:58:07-0500
This is a binomial distribution problem with n=15 and probability, p=4000/10000=0.4
Probability that at most 7 favor the new tax=p (0≤x≤7)
Using excel, the formula for binomial distribution is “=BINOM.DIST (number_s, trials, probability, cumulative)”
In this case;
Number_s=7
Trials=15
Probability=0.4
Cumulative-TRUE
Therefore, p (0≤x≤7)= BINOM.DIST(7,15,0.4,TRUE)=0.786897